# Clustering Properties of Self-Supervised Learning

### v1

@@ 5.3 Transfer to Downstream Tasks
To evaluate the quality of representations learned by ReSA, we transfer our pretrained model to downstream tasks, including COCO (Lin et al., 2014) object detection and instance segmentation. For these tasks, we adopt the baseline codebase from MoCo (He et al., 2020). Most results reported in Table 5 are inherited from SimSiam paper (Chen & He, 2021). Notably, ReSA also achieves better performance compared to other methods on both tasks, highlighting its strong potential for transfer learning in downstream applications.
@@ Acknowledgment
This work was supported by the National Science and Technology Major Project (2022ZD0116306), National Natural Science Foundation of China (Grant No. 62476016), the Fundamental Research Funds for the Central Universities.
@@ B.2 Implementation Details of ReSA Pretraining
When pretraining with Vision Transformers (ViTs), we lack the computational resources necessary for extensive hyperparameter tuning of ReSA. ViTs involve numerous intricate settings, such as initialization methods and optimization parameters, which have a significant impact on training outcomes. Notably, representative SSL methods on ViTs, such as DINO and MoCoV3, adopt totally distinct designs in both architecture and training strategies. DINO (Caron et al., 2021) leverages a range of training tricks, including weight decay scheduling, gradient clipping, and stochastic depth, among others. To stabilize training, it avoids mixed-precision training, which substantially reduces computational efficiency and increases memory requirements. MoCoV3 (Chen et al., 2021), on the other hand, proposes freezing the patch embedding layer to enhance training stability while enabling mixed-precision training. However, it still incorporates batch normalization (BN) in the projector and predictor MLPs, which reduces training efficiency and complicates its application in multi-view scenarios (Morningstar et al., 2024). Taking these considerations into account, we adopt the ViT design and initialization approach of MoCoV3 for ReSA, but remove the BN layers from the MLPs. This ensures that ReSA functions as a BN-free system during ViT training, eliminating the need for BN synchronization and improving overall training efficiency. In this paper, we set the optimizer-related parameters as shown in Table 8. We believe that further exploration of more suitable initialization methods and related parameters for ReSA could enhance its performance, as evidenced by its outstanding results on ConvNets.
@@ C.1 Ablation Studies
Meanwhile, in contrast to contrastive learning methods and other approaches such as SwAV and DINO, which typically require a small temperature value (e.g. $\tau=0.1$), ReSA achieves favorable performance with a higher temperature value ($\tau=0.4$). This indicates that the optimization process of ReSA can effectively incorporate a broader range of samples, rather than focusing exclusively on hard samples (Wang & Liu, 2021), as is the case with other methods.
Additionally, we conduct experiments to test the extraction of clustering information from embedding to obtain the soft assignment $\mathbf{A}_{\mathbf{H}}$. The results shown in Figure 10 indicate that, under this condition, the loss struggles to converge, and the model’s accuracy significantly declines compared to ReSA. This finding is consistent with the analyses in Section 3, where we note that the clustering properties of embedding are less stable and inferior to those of encoding, making it challenging for the model to effectively learn high-quality clustering information.
@@
Figure 10: Ablation on extraction of clustering information from encoding vs. embedding to obtain the soft assignment $\mathbf{A}_{\mathbf{H}}$. Both are pretrained for 1000 epochs on CIFAR-100 using ResNet-18 under totally the same experimental settings provided in Appendix B.
@@
Figure 11: Visualization of the soft assignment matrix $\mathbf{A}_{\mathbf{H}}$ during the early stages of training.
C.2 How ReSA avoids representation collapse?
@@
Although ReSA has demonstrated good performance in various experiments, it remains unclear how the model avoids feature collapse in the early stages of training when it has not yet learned any clustering information. To preliminarily address this question, we visualize the soft assignment matrix $\mathbf{A}_{\mathbf{H}}$ during the early stages of training. As shown in Figure 11, at model initialization, the assignment already exhibits a dominance of diagonal elements, reflecting an optimization trend that pulls positive samples closer together. This behavior is primarily driven by the sharp distribution induced by the small regularization parameter in the Sinkhorn-Knopp algorithm, which also contributes to pushing negative samples further apart. As training progresses, we observe that the dominance of diagonal elements gradually strengthens, effectively preventing feature collapse in the model.
C.3 The impact of weak augmentation on other methods
@@
C.4 Discussion and Future Work

### v2

@@ 5.3 Transfer to Downstream Tasks
To evaluate the quality of representations learned by ReSA, we transfer our pretrained model to downstream tasks, including COCO (Lin et al., 2014) object detection and instance segmentation. For these tasks, we adopt the baseline codebase from MoCo (He et al., 2020). Most results reported in Table 5 are inherited from SimSiam paper (Chen & He, 2021). Notably, ReSA also achieves better performance compared to other methods on both tasks, highlighting its strong potential for downstream applications.
@@ Acknowledgment
This work was supported by the National Science and Technology Major Project (2022ZD0116306), National Natural Science Foundation of China (Grant No. 62476016 and 62441617), the Fundamental Research Funds for the Central Universities.
@@ B.2 Implementation Details of ReSA Pretraining
Vision Transformer (ViT) pre-training involves numerous intricate settings, such as initialization methods and optimization parameters, which have a significant impact on training outcomes. Notably, representative SSL methods on ViTs, such as DINO and MoCoV3, adopt totally distinct designs in both architecture and training strategies. DINO (Caron et al., 2021) leverages a range of training tricks, including weight decay scheduling, gradient clipping, and stochastic depth, among others. To stabilize training, it avoids mixed-precision training, which substantially reduces computational efficiency and increases memory requirements. MoCoV3 (Chen et al., 2021), on the other hand, proposes freezing the patch embedding layer to enhance training stability while enabling mixed-precision training. However, it still incorporates batch normalization (BN) in the projector and predictor MLPs, which reduces training efficiency and complicates its application in multi-view scenarios (Morningstar et al., 2024). Taking these considerations into account, we adopt the ViT design and initialization approach of MoCoV3 for ReSA, but remove the BN layers from the MLPs. This ensures that ReSA functions as a BN-free system during ViT training, eliminating the need for BN synchronization and improving overall training efficiency. In this paper, we set the optimizer-related parameters as shown in Table 8. We believe that further exploration of more suitable initialization methods and related parameters for ReSA could enhance its performance, as evidenced by its outstanding results on ConvNets.
@@ C.1 Ablation Studies
Meanwhile, in contrast to contrastive learning methods and other approaches such as SwAV and DINO, which typically require a small temperature value (e.g. $\tau=0.1$), ReSA achieves favorable performance with a higher temperature value ($\tau=0.4$). This indicates that the optimization process of ReSA can effectively incorporate a broader range of samples with better tolerance, rather than focusing exclusively on hard samples (Wang & Liu, 2021), as is the case with other methods.
Additionally, we conduct experiments to test the extraction of clustering information from embedding to obtain the self assignment $\mathbf{A}_{\mathbf{H}}$. The results shown in Figure 10 indicate that, under this condition, the loss struggles to converge, and the model’s accuracy significantly declines compared to ReSA. This finding is consistent with the analyses in Section 3, where we note that the clustering properties of embedding are less stable and inferior to those of encoding, making it challenging for the model to effectively learn high-quality clustering information.
@@
Figure 10: Ablation on extraction of clustering information from encoding vs. embedding to obtain the self-assignment $\mathbf{A}_{\mathbf{H}}$. Both are pretrained for 1000 epochs on CIFAR-100 using ResNet-18 under totally the same experimental settings provided in Appendix B.
@@
Figure 11: Visualization of the self-assignment matrix $\mathbf{A}_{\mathbf{H}}$ during the early stages of training.
C.2 How ReSA Avoids Feature Collapse and Early Clustering Error?
Although ReSA has demonstrated good performance in various experiments, it remains unclear how the model avoids feature collapse and clustering error in the early stages of training when it has not yet learned any clustering information.
To answer these questions, we first consider the loss formula of ReSA: $\ell_{\text{ReSA}}(\mathbf{z}_{i})=-\sum_{j=1}^{m}{\mathbf{A}_{\mathbf{H}}}^{(i,j)}\log\left(\frac{\exp(s_{i,j}/\tau)}{\sum_{k=1}^{m}\exp(s_{i,k}/\tau)}\right)$. Here, ${\mathbf{A}_{\mathbf{H}}}^{(i,j)}$ denotes the $(i,j)$-th element of $\mathbf{A}_{\mathbf{H}}$, where $\mathbf{A}_{\mathbf{H}}=\text{Sinkhorn}(\mathbf{H}^{\top}\mathbf{H})$. Given that the Sinkhorn-Knopp algorithm exhibits strict monotonicity, we have the property that if ${\mathbf{h}_{i}}^{\top}{\mathbf{h}_{j}}>{\mathbf{h}_{i}}^{\top}{\mathbf{h}_{k}}$, then ${\mathbf{A}_{\mathbf{H}}}^{(i,j)}>{\mathbf{A}_{\mathbf{H}}}^{(i,k)}$. Since the vectors in $\mathbf{H}$ are $L_{2}$-normalized, the diagonal elements of $\mathbf{H}^{\top}\mathbf{H}$ are all maximized to 1. This implies that for any $i$ and $j$, we have ${\mathbf{A}_{\mathbf{H}}}^{(i,i)}\geq{\mathbf{A}_{\mathbf{H}}}^{(i,j)}$. Furthermore, due to the sharp distribution employed in Sinkhorn, the diagonal elements are significantly larger than the off-diagonal elements. This ensures that the optimization focus of ReSA remains on the alignment of augmented views from the same image, substantially reducing the impact of early assignment errors (off-diagonal elements) on the training process. Throughout training, the continual alignment enables the model to learn meaningful representations, which in turn facilitates correct cluster assignments in the later stages, further promoting learning. In contrast, other clustering-based methods such as SwAV and DINO require an initialized prototype for cluster assignment, making it more challenging to avoid early clustering errors. We speculate that this is one of the key reasons why ReSA outperforms these methods in terms of overall effectiveness.
@@ C.2 How ReSA Avoids Feature Collapse and Early Clustering Error?
We then visualize the self-assignment matrix $\mathbf{A}_{\mathbf{H}}$ during the early stages of training. As shown in Figure 11, at model initialization, the assignment already exhibits a dominance of diagonal elements, reflecting an optimization trend that pulls augmented views from the same image closer together. As training progresses, we observe that the dominance of diagonal elements gradually strengthens, effectively preventing feature collapse in the model.
C.3 Pretraining on Long-tailed Dataset
To further evaluate the training performance of ReSA on imbalanced datasets, we conduct experiments using four self-supervised learning methods—ReSA, MoCoV3, INTL, and VICReg—pretrained and evaluated on the long-tailed CIFAR100-LT dataset. We follow the setup on https://huggingface.co/datasets/tomas-gajarsky/cifar100-lt, setting an imbalance factor of 1/20, resulting in the CIFAR100-LT dataset containing 15,907 images. We train these four SSL models for 1000 epochs with ResNet-18 as the encoder and evaluate it on the full CIFAR100 test set. As observed in Figure 12, the loss values of these four methods converge well, but the final evaluation accuracy is significantly lower compared to training on the full CIFAR100 dataset in Table 1. Interestingly, we notice that ReSA’s loss decreases more slowly in the early stages, and its accuracy improves more gradually than other methods. We hypothesize that this may be due to noisy initial clusters in the early stages of training, causing clustering errors. However, we find that as training progresses, ReSA’s accuracy continues to rise in the mid-phase, surpassing all other methods. This suggests that ReSA is able to gradually resolve these issues and learn the correct clustering patterns as training advances, rather than amplifying errors. Overall, this experiment demonstrates that ReSA can also learn effective representations on long-tailed datasets.
Figure 12: Pretraining on the long-tailed dataset. Here We report the training losses of four self-supervised learning methods on the CIFAR100-LT dataset, along with their evalutaion performance on the full CIFAR-100 test set as measured by a linear probe and a k-NN classifier.
C.4 The impact of weak augmentation on other methods
@@
C.5 Discussion and Future Work